When a nylon cloth is rubbed with eyeglasses, the eyeglasses acquire a charge of 10 nC.

How many electrons were transferred during the process?



1
Expert's answer
2022-04-12T16:11:18-0400
"N=\frac{q}{e}=\frac{10*10^{-9}}{1.6*10^{-19}}=6.25*10^{10}"